# Multi-Objective Quality-Diversity for Crystal Structure Prediction

### v1

@@ 1. Introduction
Recently, Wolinska et al. (Wolinska et al., 2024) applied the well-established map-elites algorithm to the CSP problem by using domain-specific mutation operators and graph neural-networks as surrogate models for evaluating potential crystal structures. The authors demonstrate that, by using this approach, qd algorithms are capable of discovering a wide variety of crystal structures (Wolinska et al., 2024). Excitingly, their method was not only able to discover structures known by material scientists, but also to discover some promising new structures.
@@ 2.2. Quality-Diversity Algorithms
In the family of qd optimisation methods, Multi-dimensional Archive of Phenotypic Elites (map-elites) is a simple yet powerful algorithm used successfully in several prior works (Pierrot et al., 2022; Nickelson et al., 2023; Janmohamed et al., 2023). In map-elites, the feature space is discretised into a grid composed of cells. Each cell corresponds to a feature vector and stores the solution characterised by the same feature. The algorithm begins with initialising the grid and placing randomly sampled solutions in the cells. At each iteration of the map-elites algorithm, the first step consists of selecting a batch of solutions from the grid. Next, copies of these solutions undergo variation to form offspring solutions, which are evaluated to obtain the fitness scores and the features. Finally, each new solution is added back to the grid if the cell matching the solution’s feature vector is empty or if the existing solution within the cell is outperformed by the new solution. By repeating this loop for a specified budget, the grid gradually accumulates diverse and high-quality solutions. The performance of map-elites algorithms can be assessed via the qd-score which is the sum of the fitness of the solutions in each cell.
@@ 2.4. Multi-Objective Quality-Diversity
Multi-Objective Quality-Diversity (moqd) aims to generate a diverse set of high-performing solutions that maximise a combination of $k$ different objectives $\textbf{f}=[f_{1},f_{2},...f_{k}]$. Multi-Objective MAP-Elites (mome) (Pierrot et al., 2022), a well-established moqd algorithm, achieves this by storing a Pareto Front of solutions in each grid cell. The algorithm flow closely matches that of MAP-Elites with one crucial distinction: after evaluating a new solution and identifying its corresponding cell based on its feature, the solution is placed into the cell only if it Pareto-dominates the other existing solutions in that cell. The objective of moqd algorithms is to find an approximation of the Pareto Front in each cell that has the best set of trade-offs and therefore maximum hypervolume. Hence, the performance of moqd algorithms is assessed by the sum of the hypervolumes of all the Pareto Front approximations in the grid, referred to as the moqd-score. This objective translates to finding the best possible trade-offs across the set of objectives for each feature of interest.
@@ 5.2. Baselines
We evaluate our method, mome-x, against three map-elites baselines: 1) me-energy 2) me-magnetism and 3) me-sum. Each of these baselines corresponds to applying the map-elites algorithm to CSP, optimising for stability of structures, magnetism of structures and an additive sum of the two respectively. We use the same mutation operators, initialisation and relaxation procedures for all of the baselines and mome-x. However, it should be noted that a straightforward comparison between mome algorithms and map-elites algorithms is non-trivial, since map-elites algorithms store at most one solution per cell whereas mome algorithms can store several. In particular, if we use an archive tessellated into $c$ cells for mome-x, each with a maximum Pareto Front length $p$, mome-x could have a maximum population size of $c\times p$. Therefore, for fairness, in all map-elites baselines we tessellate the archive into $c\times p$ cells.
@@ 6.1. Multi-Objective Quality-Diversity Results
    energy-qd-score: the sum of the energy scores of all solutions in the archive. This is what the me-energy baseline is explicitly trying to maximise.
@@
Figure 4 shows also that mome-x also outperforms ($p<0.03$) all other baselines on the energy-qd-score and magnetism-qd-score. This is a particularly intriguing result, given that these metrics are what me-energy and me-magnetism are respectively explicitly aiming to maximise. This corroborates previous observations that simultaneously optimising over several objectives may provide helpful stepping stone solutions that aid exploration and, in turn, can help to maximise each objective separately (Janmohamed et al., 2023). Additionally, we observe that mome-x also outperforms me-sum on these metrics. This could suggest that me-sum succeeds in finding solutions that are balanced across the two objectives, but is less able to find solutions that are high-performing on each objective separately and therefore provides fewer possible trade-offs for each cell.
@@ 6.3. Evaluation Against Reference Structures
mome-x me-energy me-magnetism me-sum

### v2

@@ 1. Introduction
Recently, Wolinska et al. (Wolinska et al., 2024) applied the well-established map-elites (Mouret and Clune, 2015) algorithm to the CSP problem by using domain-specific mutation operators and graph neural-networks as surrogate models for evaluating potential crystal structures. The authors demonstrate that, by using this approach, qd algorithms are capable of discovering a wide variety of crystal structures (Wolinska et al., 2024). Excitingly, their method was not only able to discover structures known by material scientists, but also to discover some promising new structures.
@@ 2.2. Quality-Diversity Algorithms
In the family of qd optimisation methods, Multi-dimensional Archive of Phenotypic Elites (map-elites) is a simple yet powerful algorithm used successfully in several prior works (Mouret and Clune, 2015; Cully et al., 2015; Cully, 2021). In map-elites, the feature space is discretised into a grid composed of cells. Each cell corresponds to a feature vector and stores the solution characterised by the same feature. The algorithm begins with initialising the grid and placing randomly sampled solutions in the cells. At each iteration of the map-elites algorithm, the first step consists of selecting a batch of solutions from the grid. Next, copies of these solutions undergo variation to form offspring solutions, which are evaluated to obtain the fitness scores and the features. Finally, each new solution is added back to the grid if the cell matching the solution’s feature vector is empty or if the existing solution within the cell is outperformed by the new solution. By repeating this loop for a specified budget, the grid gradually accumulates diverse and high-quality solutions. The performance of map-elites algorithms can be assessed via the qd-score which is the sum of the fitness of the solutions in each cell.
@@ 2.4. Multi-Objective Quality-Diversity
Multi-Objective Quality-Diversity (moqd) aims to generate a diverse set of high-performing solutions that maximise a combination of $k$ different objectives $\textbf{f}=[f_{1},f_{2},...f_{k}]$. Multi-Objective MAP-Elites (mome) (Pierrot et al., 2022), a well-established moqd algorithm, achieves this by storing a Pareto Front of solutions in each grid cell. The algorithm flow closely matches that of MAP-Elites with one crucial distinction: after evaluating a new solution and identifying its corresponding cell based on its feature, the solution is placed into the cell only if it Pareto-dominates the other existing solutions in that cell. The objective of moqd algorithms is to find an approximation of the Pareto Front in each cell that has the best set of trade-offs and therefore maximum hypervolume (Pierrot et al., 2022; Nickelson et al., 2023; Janmohamed et al., 2023). Hence, the performance of moqd algorithms is assessed by the sum of the hypervolumes of all the Pareto Front approximations in the grid, referred to as the moqd-score. This objective translates to finding the best possible trade-offs across the set of objectives for each feature of interest.
@@ 5.2. Baselines
We evaluate our method, mome-x, against three map-elites baselines: 1) me-stability 2) me-magnetism and 3) me-sum. Each of these baselines corresponds to applying the map-elites algorithm to CSP, optimising for stability, magnetism and an additive sum of the two respectively. We use the same mutation operators, initialisation and relaxation procedures for all of the baselines and mome-x. However, it should be noted that a straightforward comparison between mome algorithms and map-elites algorithms is non-trivial, since map-elites algorithms store at most one solution per cell whereas mome algorithms can store several. In particular, if we use an archive tessellated into $c$ cells for mome-x, each with a maximum Pareto Front length $p$, mome-x could have a maximum population size of $c\times p$. Therefore, for fairness, in all map-elites baselines we tessellate the archive into $c\times p$ cells.
@@ 6.1. Multi-Objective Quality-Diversity Results
    energy-qd-score: the sum of the energy scores of all solutions in the archive. This is what the me-stability baseline is explicitly trying to maximise.
@@
Figure 4 shows also that mome-x also outperforms ($p<0.03$) all other baselines on the energy-qd-score and magnetism-qd-score. This is a particularly intriguing result, given that these metrics are what me-stability and me-magnetism are respectively explicitly aiming to maximise. This corroborates previous observations that simultaneously optimising over several objectives may provide helpful stepping stone solutions that aid exploration and, in turn, can help to maximise each objective separately (Janmohamed et al., 2023). Additionally, we observe that mome-x also outperforms me-sum on these metrics. This could suggest that me-sum succeeds in finding solutions that are balanced across the two objectives, but is less able to find solutions that are high-performing on each objective separately and therefore provides fewer possible trade-offs for each cell.
@@ 6.3. Evaluation Against Reference Structures
mome-x me-stability me-magnetism me-sum